OpenStudy (anonymous):

In the triangle shown below, find the measurement of Angle B. Round to the nearer degree.

OpenStudy (anonymous):



OpenStudy (anonymous):

It would be great someone knew the answer. i'm kind of in a hurry...

OpenStudy (anonymous):

the tan B = 8/15 so find the atan (8/15)

OpenStudy (anonymous):

It didn't come out right..?

OpenStudy (anonymous):

\[\tan^{-1} \left( \frac{ 8 }{ 15 } \right)\]

OpenStudy (anonymous):

find this and you find the angle you need

OpenStudy (anonymous):

do you have a calculator on your computer?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

use it

OpenStudy (anonymous):

you may have to put it into scientific mode

OpenStudy (anonymous):

I did , but it came out as ).57 or something.. and I was in degree mode.. :(

OpenStudy (anonymous):

that look more like radians... on mine, I had to first divide 8 by 15 which gave .5333333333 and then push the inverse button and then the tan button

OpenStudy (anonymous):

let me try again :)

OpenStudy (anonymous):

I still got .5-

OpenStudy (anonymous):

switch to degree mode.

OpenStudy (anonymous):

I got 0.009...and i'm pretty sure that's nor right...

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

Can someone please help me i'm kind of in a rush.. my options are A.28 B.30 C.32 D.34

OpenStudy (anonymous):

28

OpenStudy (anonymous):

Thankyou! Can you help me with one more similar to this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

start a new question tho

OpenStudy (anonymous):

In Triangle ABC shown below, use an appropriate trigonometric ratio to find measure B. round to the nearer degree.